OpenStudy (anonymous):

what is the simplified form of (3x+3/x-2)-(x-4/x-2)

OpenStudy (anonymous):

helpp /:

hero (hero):

\(\large\frac{3x+3}{x-2} - \frac{x-4}{x-2}\)

OpenStudy (anonymous):

what do i do with this ?

hero (hero):

Combine fractions to get: \(\large\frac{3x + 3 - (x-4)}{x-2}\)

OpenStudy (anonymous):

then end ?

hero (hero):

Nope, you have to simplify the numerator.

OpenStudy (anonymous):

okay .

OpenStudy (anonymous):

i got 3x^2-9x-12

hero (hero):

I have no idea how you got that. (3x+3) - (x-4) are two expressions that you subtract. (3x+3)(x-4) is a multiplication, but that's not the same as the above expression.

OpenStudy (anonymous):

2x+7?

hero (hero):

Yes that would be the numerator

hero (hero):

So what would the full simplified fraction look like?

OpenStudy (anonymous):

2x+7/x-2?